HUBBLE SPACE TELESCOPE DAILY REPORT #2424 PERIOD COVERED: 0000Z (UTC) 07/28/99 - 0000Z (UTC) 07/29/99 Daily Status Report as of 210/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.2 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed WF/PC-2 8195 (The Stellar Halo and the Metallicity Distribution Function in the Giant Elliptical NGC 5128) The WF/PC-2 was used to make observations of the giant elliptical galaxy NGC-5128 in order to study its stellar halo and the metallicity distribution function. Because this galaxy is 2 magnitudes closer than any other giant elliptical galaxy, it gives us a unique opportunity to study the enrichment history at a level of detail that is impossible for any other such galaxy. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of WF/PC-2 8230 (Snapshot Survey of Dynamically Close Galaxy Pairs from z=0.1 to z=0.5) The WF/PC-2 was used to make snapshot survey observations of galactic cluster Gal-Clus-2723. This is to obtain snapshot images of a large, well-defined, and dynamically selected sample of close galaxy pairs at moderate redshift. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Three Sets of WF/PC-2 8053 (WF/PC-2 Cycle-7 Earth Flats -- Part II) The WF/PC-2 was used to take Earth flat calibration images. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 8192 (A Snapshot Survey of Probable Nearby Galaxies) The WF/PC-2 was to be used to make a snapshot survey images of a probable nearby galaxy KK009. As noted in HSTAR 7203 and 2.1, the acquisition for this proposal failed and the take data flag remained down. 1.8 Completed WF/PC-2 8450 (WF/PC-2 Cycle-8 Noiseless Preflash Test) The WF/PC-2 was used to make observations of the large globular cluster Omega-Cen, as part of a Cycle-8 noiseless preflash test. This program will test whether a noiseless preflash can reduce CTE and long vs. short effects. An internal flat and a dark calibration image were also taken. During this proposal, the re-acquisition failure described in HSTARs 7204 and 7205, as well as 2.1, occurred, and the take data flag was down for a portion of the observations. No further anomalies were noted. 1.9 Completed STIS/CCD 8410 (STIS CCD Hot Pixel Annealing C8) The Space Telescope Imaging Spectrograph (CCD) was used to assess the effectiveness of the CCD hot pixel annealing process by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. In addition CTE performance is examined by looking for traps in a low signal level flat. During this proposal, the re-acquisition failure described in HSTARs 7204 and 7205, as well as 2.1, occurred, and the take data flag was down for a portion of the observations. No further anomalies were noted. 1.10 Completed STIS/MA1 8316 (The Lyman-Alpha Absorption Cross-Section of Nearby Galaxies) > The Space Telescope Imaging Spectrograph (MA1) was used to study the Lyman-Alpha absorption cross-section of nearby galaxy TON730. The high resolution of the spectra will allow us to accurately determine the velocity of the gas, which can be compared to the systemic velocity of the galaxy, and to the velocity field determined from 21 cm H-I emission measurements, when available, enabling us to determine if or how the gas is kinematically tied to the galaxy. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 Per HSTAR 7203, the acquisition scheduled for 209/123000Z failed when the search radius limit was exceeded. The proposal detailed in 1.7 was affected. Scheduled Reacquisitions : 8 Successful : 5 HSTAR 7204 described a saved error null bias which was out of limits at 209/150000Z. Both FGSs acquired, and the acquisition was successful. However, as a result, a following re-acquisitions at 209/1754Z, at 209/1930Z, and at 209/2106Z failed due to the scan step limit being exceeded. HSTAR 7205 was written. A real-time map was executed at 209/1911Z per an operations request. The proposals described in 1.8 and 1.9 were at least partially affected. 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: The operations team continues to execute its functions from the SAMS back-up facility. JIS-7 (EVA-1) was performed yesterday. JIS-1 (Rendezvous/Capture) will begin today and run until tomorrow morning. The SSR EDAC error counter was cleared twice using ROP SR-1. The RMGA calibration test, part 1, was completed successfully at 210/0038Z. Parts 2 and 3 were completed at 210/0245Z. An NSSC-1 status buffer dumped was performed at 210/0348Z per ROP NS-3. During an SAA passage and during an SAA, at 210/081948Z, the search radius limit flag was raised even though the FGSs were not guiding at the time. HSTAR 7206 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH